A square and a triangle have equal perimeters. The lengths of the
sides of the triangle are 6.2 cm, 8.2 cm and 9.6 cm. What is the area of
the square in centimetres?
a) 24 b) 25 c) 36 d) 48 e) 64

Answer by toidayma(44)

The perimeter of the triangle is: 6.2 + 8.2 + 9.6 = 24(cm) = the perimeter of the square, which means that a size's length of the square is: 24/4 = 6(cm). Therefore, the area of the square is 6*6 = 36(cm2) = answer choice (C).